A Cheez-It box weighs 16 ounces. The serving size is 12 crackers. There are 140 calories in each ounce. A box contains 32 servings. Calculate the number of calories per cracker...
What?!?!

Answer by hyperasad(6)

weight of cheese- it box= 16 ounces
serving size = 12 crackers
calories per ounce = 140 calories
box contains = 32 servings
weight of cheese per serving = 16/32 = 0.5 ounce
calories in each serving = 140x0.5=70 calories
(as each serving contains 12 crackers, divide 70 calories by No. of crackers in each serving)
Calories per cracker = 70/12=5.8333 calories